OpenStudy (anonymous):

What is (4 1/3) / (2 3/5) ? Express the answer in simplest form.

OpenStudy (igreen):

Convert them into improper fractions.

OpenStudy (igreen):

Do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (igreen):

Okay, tell me what you come up with.

OpenStudy (anonymous):

13/3 and 13/5

OpenStudy (igreen):

Yes, now when dividing fractions, we flip the fraction we're dividing by and multiply. \(\sf\dfrac{13}{3} \div \dfrac{13}{5} \rightarrow \dfrac{13}{3} \times \dfrac{5}{13}\)

OpenStudy (igreen):

Can you multiply those two fractions together?

OpenStudy (anonymous):

i know that way, its kcf *keep, change, flip*

OpenStudy (igreen):

Oh, for a second there I thought you meant Kentucky Fried Chicken

OpenStudy (anonymous):

13 * 5 is 65. 3 / 13 39

OpenStudy (igreen):

3 * 13 is 39, yes

OpenStudy (igreen):

So we have: \(\sf \dfrac{65}{39}\) That's our answer, unless it wants it in mixed number form.

OpenStudy (anonymous):

then now all i have to do is break it down to the simplest form?

OpenStudy (anonymous):

i broke it down, and the answer i got is 1 6/10 im going to break it down more and say 1 3/5 thats the lowest i can get, thanks for the help

OpenStudy (igreen):

Not quite

OpenStudy (igreen):

1 3/5 is incorrect

OpenStudy (anonymous):

? then i did 69/39 it gave me 1.666 then i rounded it to 1 and 6/10 it asked to be in simplest form,so then i divided by 2, for 1 and 3/5, i cant break it down anymore, so what do i do now?